OpenStudy (henrietepurina):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10

OpenStudy (henrietepurina):

-2y + 10 2y + 10 2y - 10 -2y - 10

OpenStudy (henrietepurina):

@ParthKohli

OpenStudy (sleepyhead314):

try to solve for x from -x - 2y = -10

OpenStudy (sleepyhead314):

to do that you can add 2y to both sides then divide both sides by a -1

OpenStudy (henrietepurina):

3x + 2y = -8 x = -10 am I close?

OpenStudy (sleepyhead314):

hmmm not quite all they want is for you to isolate x from the equation of -x - 2y = -10 so add 2y to both sides +2y +2y so -x = 2y - 10 then divide both sides by -1

OpenStudy (henrietepurina):

is it B cuz x = 2y + 10

OpenStudy (sleepyhead314):

ooooooo so close so close ^_^ don't forget that multiplying/dividing by -1 would mean that -(2y-10) and distribute the negative

OpenStudy (henrietepurina):

-2y + 10?

OpenStudy (sleepyhead314):

correct ^_^

OpenStudy (henrietepurina):

yay thanks :)

OpenStudy (sleepyhead314):

you're welcome :)